OpenStudy (asnaseer):

When typing up a reply, is it possible to enter equations without forcing them to start on the next line? i.e. I want to be able to type something like: "some random reply \[\frac{3}{4}\] is shown here" with the fraction 3/4 appearing inline with the text.

OpenStudy (asnaseer):

so the reply should end up something like this: "some random reply 3/4 is shown here"

OpenStudy (anonymous):

I think the openstudy team should enable this option because I have also had such problems(not being able to type the equation in the same line).

OpenStudy (asnaseer):

It must be possible because I have seem some people type their replies exactly I want. I just don't know how they did it.

OpenStudy (asnaseer):

** exactly as I ...

OpenStudy (anonymous):

but I haven't seen any such replies.

OpenStudy (anonymous):

yes it's possible look \(\frac{1}{2}\) is on the same line!

OpenStudy (anonymous):

you need to write \\( \\) just single slashes

OpenStudy (asnaseer):

let me try \(\frac{3}{4}\) - wow! thanks @Thomas

OpenStudy (asnaseer):

After more investigation I have also found that in cases where you need a little bit of text in what is mostly an equation you can use \mbox{some text}. For example: \[\mbox{using }x=\frac{-b\pm\sqrt{b^2-4ac}}{2a}\mbox{ leads to }x=1\mbox{ or }2\]

OpenStudy (anonymous):

\[\mbox{a}x^2+\mbox{b}x+\mbox{c}=0\] only x should be emphasized because it's variable :P

OpenStudy (shadowfiend):

First off, you can also use \text{}, which is easier to remember. Secondly, a, b, and c should all be italicized in the above. They are all placeholders, despite the fact that they are constants. Only straight text should be non-italicized.